Could you help me with problem. Is it possible to explain step by step.Thank You
When a = 0.1, b = 0.5 and c= 0.2
a-(b-c)=0.1 - (.05 - .02) = ?

Answer by jaydducote(11)

When a = 0.1, b = 0.5 and c= 0.2
a-(b-c)=0.1 - (.05 - .02) = ?
This is right so far. Remember the order of operations PEMDAS, so you have to do what is in the parentheses first.
(.05-.02) = .03
so now you have .01-.03=?
.01 is smaller than .03, so you will get a negative answer.
anytime you are adding or subtracting positives and negatives, if you have the same sign, add and keep the sign. If you have opposite signs, subtract and keep the sign of the larger number. so with .01-.03, you have opposite signs. Subtract .03-.01=.02 However, since .03 is larger than .01, you'll need to take that sign, which is negative. The answer is -.02